Calculate the potential of single electrode
a. Zn2+(aq) (0.1M) / Zn. When Eo= -0.76V.
b. Ag+ (0.01M) / Ag (s) E0 = 0.799
1
Expert's answer
2020-05-10T09:19:35-0400

At anode

"Zn\to Zn^{2+}+2e^-" ".....E_0=0.76 V.................(1)"


at cathode

"Ag^++e^-\to Ag" ........."E_0=0.799................(2)"


(1)+2"\times"(2)

"Zn+2Ag^+\to Zn^{2+}+2Ag"

"E_{cell}=0.76V +0.799V=1.559 V"